OpenStudy (anonymous):

Write the partial fraction decompression for the rational expression (2X + 5)/(x^2 -x -2)

OpenStudy (anonymous):

does the denominator factor?

OpenStudy (anonymous):

\[\frac{2x+5}{(x^2-x-2)}=\frac{2x+5}{(x-2)(x+1)}=\frac{A}{x-2}+\frac{B}{x+1}\] i think

OpenStudy (anonymous):

you want a quick way to do it?

OpenStudy (anonymous):

ya sure

OpenStudy (anonymous):

ok lets find A what makes \(x-2=0\)?

OpenStudy (anonymous):

if x was 2

OpenStudy (anonymous):

right now take this \[\frac{2x+5}{(x-2)(x+1)}\] and put your hand over that factor \[\frac{2x+5}{\cancel{(x-2)}(x+1)}\] then replace \(x\) by \(2\) and get \[\frac{2\times 2+5}{2+1}=\frac{9}{3}=3\] and so \(A=3\)

OpenStudy (anonymous):

oh I get it so B=-1?

OpenStudy (anonymous):

for the next one do the same thing, this time use \(x=-1\) and put your hand over the other factor \[\frac{2x+5}{(x-2)\cancel{(x+1)}}\]

OpenStudy (anonymous):

yeah that is it

OpenStudy (anonymous):

quick right?

OpenStudy (anonymous):

yu

OpenStudy (anonymous):

otherwise you write \[A(x+1)+B(x-2)=2x+5\] and solve for \(A\) and \(B\) but it amounts to the same thing

OpenStudy (anonymous):

oops yup. Thanks a lot for your help!

OpenStudy (anonymous):

yw